OpenStudy (anonymous):

Solve this system of equations by using substitution. 3x + y = -6 y = x + 10 The answer (x,y)

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

they tell you what y is equal to ... so substitute that into all y parts

OpenStudy (amistre64):

y = x + 10 3x + y = -6 y = x + 10 3x + (x + 10) = -6 x + 10 = x + 10 the second equation is always true so the only useful one we can use is the first equation ... solve for x when we know what the number value of x is, we still know that y = x+10